### v4

# A Survey on Vision-Language-Action Models for Embodied AI

Abstract
Embodied AI is widely recognized as a key element of artificial general intelligence because it involves controlling embodied agents to perform tasks in the physical world. Building on the success of large language models and vision-language models, a new category of multimodal models—referred to as vision-language-action models (VLAs)—has emerged to address language-conditioned robotic tasks in embodied AI by leveraging their distinct ability to generate actions. In recent years, a myriad of VLAs have been developed, making it imperative to capture the rapidly evolving landscape through a comprehensive survey. To this end, we present the first survey on VLAs for embodied AI. This work provides a detailed taxonomy of VLAs, organized into three major lines of research. The first line focuses on individual components of VLAs. The second line is dedicated to developing control policies adept at predicting low-level actions. The third line comprises high-level task planners capable of decomposing long-horizon tasks into a sequence of subtasks, thereby guiding VLAs to follow more general user instructions. Furthermore, we provide an extensive summary of relevant resources, including datasets, simulators, and benchmarks. Finally, we discuss the challenges faced by VLAs and outline promising future directions in embodied AI. We have created a project associated with this survey, which is available at https://github.com/yueen-ma/Awesome-VLA.
Index Terms:
Vision-Language-Action Model, Embodied AI, Robotics, Multi-modality, Large Language Model, World Model
I Introduction
Vision-language-action models (VLAs) are a class of multimodal models within the field of embodied AI, designed to process information from vision, language, and action modalities. Unlike conversational AI such as ChatGPT [1], embodied AI requires controlling physical embodiments that interact with the environment and robotics is the most prominent domain of embodied AI. In language-conditioned robotic tasks, the policy must possess the capability to understand language instructions, visually perceive the environment, and generate appropriate actions, necessitating the multimodal capabilities of VLAs. The term was recently coined by RT-2 [2]. Compared to earlier deep reinforcement learning approaches, VLAs offer superior versatility, dexterity, and generalizability in complex environments. As a result, they are well-suited not only for controlled settings like factories but also for everyday tasks in household environments.
Early developments in deep learning primarily consist of unimodal models. In computer vision (CV), models like AlexNet [3] showcased the potential of artifical neural networks. Recurrent neural networks laid the groundwork for numerous natural language processing (NLP) models but have seen a transition in recent years, with Transformers [4] taking precedence. The Deep Q-network [5] demonstrated that ANNs can successfully tackle reinforcement learning problems. Leveraging advancements of unimodal models across diverse machine learning fields, multimodal models now address a wide range of tasks, such as visual question answering, image captioning, and text-to-video generation.
Conventional robot policies based on reinforcement learning have largely focused on addressing a limited set of tasks within controlled environments, such as item grasping [6]. However, there is a growing need for more versatile multi-task policies, akin to recent trends in large language models (LLMs) and vision-language models (VLMs). Developing a multi-task policy is challenging, as it requires learning a broader set of skills and adapting to diverse environments. Furthermore, specifying tasks via language instructions offers a more intuitive user-robot interface, which necessitates the development of language-conditioned robot policies.
Figure 1: The general architecture of vision-language-action models. Important related components are shown in the dashed-line boxes.
Built upon the success of large VLMs, vision-language-action models have demonstrated their potential in addressing these challenges, as illustrated in Figure 1. Similar to VLMs, VLAs utilize vision foundation models as vision encoders to obtain pretrained visual representations of the current environmental state, such as object class, pose, and geometry. VLAs encode instructions using the token embeddings of their LLMs and employ various strategies to align vision and language embeddings, including approaches like BLIP-2 [7] and LLaVA [8]. By finetuning on robot data, the LLM can function as a decoder to predict actions and perform language-conditioned robotic tasks. These cross-disciplinary innovations drive the advancement of VLAs in embodied AI—a critical building block for artificial general intelligence (AGI).
(a) Concepts
(b) Timelines
Figure 2: (a) A Venn diagram that outlines the main concepts in embodied AI discussed in this paper. (b) Timelines that trace the evolution from unimodal models to vision-language-action models.
VLAs are closely related to three lines of work, as depicted by the timelines in Figure 2b and the taxonomy in Figure 3. Some approaches focus on individual components of VLAs (§III-A), such as pretrained visual representations, dynamics learning, world models, and reasoning. Meanwhile, a substantial body of research is dedicated to low-level control policies (§III-B). In this category, language instructions and visual perceptions are fed into the control policy, which then generates low-level actions—such as translation and rotation—thereby rendering VLAs an ideal choice for control policies. In contrast, another category of models serves as high-level task planners responsible for task decomposition (§IV). These models break down long-horizon tasks into a sequence of subtasks that, in turn, guide VLAs toward the overall goal, as illustrated in Figure 4. Most current robotic systems adopt such a hierarchical framework [9, 10, 11] because the high-level task planner can leverage models with high capacity while the low-level control policy can focus on speed and precision, similar to hierarchical reinforcement learning.
To provide a more comprehensive overview of current progress in embodied AI, we propose a generalized definition of “VLA,” as illustrated by the Venn diagram in Figure 2a. We define a VLA as any model capable of processing multimodal inputs from vision and language to produce robot actions that accomplish embodied tasks, typically following the architecture in Figure 1. The original concept of VLA referred to a model that adapts VLMs to robotic tasks [2]. Analogous to the distinction between LLMs and more general language models, we designate the original VLAs as “Large VLAs” (LVLAs) because they are based on LLMs or large VLMs.
Related Work. To the best of our knowledge, this survey is the first to review the recent progress of VLA models, a rapidly emerging research area. Previous surveys have investigated other facets of embodied AI. [12] comprehensively summarized foundation models in robotics up to 2023, while [13] focused on LLMs in robotics. [14] examined more recent vision, language, and robotic foundation models for general-purpose robots. [15] concentrates on real-world robot applications. In contrast, our work emphasizes VLA models, thereby complementing and extending the existing literature on embodied AI.
Contributions. To the best of our knowledge, this paper is the first comprehensive survey of emerging vision-language-action (VLA) models in the field of embodied AI.
  • Comprehensive Review. We present a thorough review of emerging VLA models in embodied AI, covering various aspects including components, architectures, training objectives, robotic tasks, etc.
  • Taxonomy. We introduce a taxonomy of VLAs based on the hierarchical framework of current robotic systems, comprising two hierarchies: a low-level control policy and a high-level task planner. Control policies execute low-level actions based on specified language commands and the perceived environment. Task planners provide guidance to control policies by decomposing long-horizon tasks into subtasks. We also discuss various essential components of VLAs.
  • Resources. We summarize the necessary resources for training and evaluating VLA models, including recent datasets and benchmarks in real-world or simulated environments. We also discuss various approaches to address current challenges, such as data scarcity and inconsistency.
  • Future Directions. We outline current challenges and promising future opportunities in the field, such as safety, foundation models, and real-world deployment.
II Background
[…]
Figure 3: The taxonomy of VLA models. The organization of this survey follows this taxonomy.
II-A Unimodal Models
Vision-language-action models integrate three modalities, often relying on existing unimodal models. The transition from convolutional neural networks (e.g., ResNet [16]) to visual Transformers (e.g., ViT [17], SAM [18]) in computer vision has facilitated the development of vision foundation models (VFMs). In natural language processing, the evolution from recurrent neural networks (e.g., LSTM [19], GRU [20]) to Transformers [4] initially led to the pretrain-finetune paradigm (e.g., BERT [21], ChatGPT [1]), followed by the recent success of prompt tuning driven by large language models. Reinforcement learning (e.g., DQN [5], AlphaGo [22], PPO [23], Dactyl [24]) has also witnessed a shift towards employing Transformers (§III-A1) to model the Markov Decision Process as autoregressive sequential data.
II-B Vision-Language Models
Vision-language tasks, encompassing image captioning [25], visual question answering [26], visual grounding [27], require the fusion of computer vision and natural language processing models. Early efforts, such as Show and Tell [28], leveraged the success of early CNNs and RNNs. The advent of high-capacity language models, including BERT [21] and GPT [29], ushered in a new era of Transformer-based VLMs. One of the pioneering self-supervised pretraining methods is ViLBERT [30], while CLIP [31] popularized contrastive pretraining approaches for multimodal alignment. The recent emergence of large language models has driven the development of multimodal LLMs (MLLMs) or large multimodal models (LMMs), which achieve state-of-the-art performance on multimodal instruction-following tasks. Representative MLLMs include Flamingo [32], BLIP-2 [7], and LLaVA [8]. VLMs share a close relationship with VLAs, as the multimodal architectures of VLMs can be readily adopted for VLAs. Additionally, VLMs can function as high-level task planners if they possess sufficient reasoning capabilities.
II-C Embodied AI & Robot Learning
Embodied AI is a unique form of artificial intelligence that actively interacts with the physical environment. This sets it apart from other AI models, such as conversational AI, which primarily handles textual conversations (ChatGPT [1]), or generative AI models that focus on tasks like text-to-video generation (Sora [33]). Embodied AI encompasses a broad spectrum of embodiments, including smart appliances, smart glasses, autonomous vehicles, and more. Among them all, robots stand out as one of the most prominent embodiments.
Robot learning is also usually framed as reinforcement learning problems, represented as a Markov Decision Process (MDP) consisting of states ($s$), actions ($a$), and rewards ($r$). A MDP trajectory is denoted as $\tau=(s_{1},a_{1},r_{1},\dots,s_{T},a_{T},r_{T})$. In certain scenarios, robotic tasks may also be viewed as partially-observable Markov Decision Processes (POMDPs) due to incomplete observations. The primary objective of reinforcement learning is to train a policy capable of generating optimal actions for the current state $\pi(a_{t}|s_{\leq t},a_{<t})$. Various methods such as TD learning, policy gradients, etc., can be employed to achieve this. However, in cases where defining the reward function proves challenging, imitation learning is utilized to directly model the action distribution within trajectories devoid of rewards $\tau=(s_{1},a_{1},\dots,s_{T},a_{T})$. Furthermore, many multi-task robot models employ language as instructions $p$ to determine which task or skill to execute, leading to the development of language-conditioned robot policies $\pi(a_{t}|p,s_{\leq t},a_{<t})$.
III Vision-Language-Action Models
III-A Components of VLA
A body of work focuses on individual components of VLA models, drawing on successes in CV, NLP, and RL. We introduce them through the lens of embodied AI.
III-A1 Reinforcement Learning
Reinforcement learning (RL) trajectories, characterized by state-action-reward sequences, naturally align with the structure of sequence modeling problems, rendering them well-suited for Transformer models. These RL Transformers have become foundational to many recent VLA models. Pioneering efforts in this direction include Decision Transformer (DT) [34] and Trajectory Transformer (TT) [35]. Gato [36] further extends this paradigm to a multi-modal/task/embodiment setting.
Reinforcement learning from human feedback (RLHF) has become an important training ingredient for LLMs. SEED [37] utilizes RLHF alongside skill-based reinforcement learning to address the sparse reward issue of long-horizon tasks in robot learning. Reflexion [38] proposes a novel verbal reinforcement learning framework, replacing weight updates in RL models with linguistic feedback.
III-A2 Pretrained Visual Representations
The effectiveness of the vision encoder directly influences the performance of VLAs, as it provides crucial information regarding the current state $s_{t}$, such as object categories, positions, and affordance. Consequently, numerous methods are devoted to improving the quality of pretrained visual representations (PVRs). Their technical details are compared in Table I.
CLIP [31] has gained widespread adoption as a vision encoder in robotic models. The training objective of CLIP is to identify the correct text-image pair among all possible combinations in a given batch. CLIP undergoes training on the WIT dataset, comprising 400 million image-text pairs. This large-scale training allows CLIP to develop a rich understanding of the relationship between visual and textual information.
R3M [39] proposes two main pretraining objectives: time contrastive learning and video-language alignment. The objective of time contrastive learning is to minimize the distance between video frames that are temporally close while simultaneously increasing the separation between frames that are temporally distant. This objective aims to create PVRs that capture the temporal relationship within the video sequence. On the other hand, video-language alignment is to learn whether a video corresponds to a language instruction. This objective enriches the semantic relevance embedded in the PVRs. VIP [40] also capitalizes on video temporal relationship.
MVP [41] adopts masked autoencoder (MAE) from computer vision to robotic datasets. The MAE involves masking out a portion of input patches to a ViT model and training it to reconstruct the corrupted patches. This approach closely resembles the masked language modeling technique used in BERT [21] and falls within the purview of self-supervised training. RPT [42] undergoes pretraining with a focus not only on reconstructing visual inputs but also on robotic actions and proprioceptive states.
Table I: Pretrained visual representations. V: Vision. L: Language. CL: contrastive learning. TFM: Transformer. Sim/Real: simulated/real-world. Mani/Navi: manipulation/navigation. For simplicity, we only show the main part of the objective equation, omitting elements such as temperature, auxiliary loss, etc. $\mathcal{S}(\cdot)$ is some similarity measurement.
Model V-Encoder Type Objective Notation Robotic Tasks
CLIP [31] ViT-B [17] VL-CL $\sum_{i=1}^{N}-\log\frac{\exp\left(\mathcal{S}(x_{i},y_{i})\right)}{\sum_{j=1}^{N}\exp\left(\mathcal{S}(x_{i},y_{j})\right)}$ $(x_{i},y_{i})$: image-text pair (Used by CLIPort [43], EmbCLIP [44], and CoW [45])
R3M [39] ResNet-50 [16] Time-CL $\sum_{i=1}^{T}-\log\frac{\exp\left(\mathcal{S}(x_{i},x_{j})\right)}{\exp\left(\mathcal{S}(x_{i},x_{j})\right)+\exp\left(\mathcal{S}(x_{i},x_{k})\right)}$ $x_{i}$: $i$-th video frame, $i<j<k$ Sim-Mani: Meta-World, Franka Kitchen, Adroit
MVP [41] ViT-B
ViT-L
MAE [46] $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i}\right)$ $x_{i}$: image patch Real-Mani (xArm7): pick, reach block, push cube, close fridge
VIP [40] ResNet-50 Time-CL $\sum_{i=1}^{T}-\log\frac{\exp\left(\mathcal{S}(x_{0},x_{j})\right)}{\exp\left(\mathcal{S}(x_{0},x_{j})\right)+\exp\left(\mathcal{S}(x_{i},x_{j})\right)}$ $x_{i}$: $i$-th video frame, $0<i<j$ Real-Mani (Franka): pick and place, close drawer, push bottle, fold towel
VC-1 [47] ViT-L MAE, CL (A study of previous works, CLIP, R3M, MVP, VIP.) Sim-Mani: Meta-World, Adroit, DMC, TriFinger, Habitat 2.0; Sim-Navi: Habitat
Voltron [48] TFM [4] MAE, Lang-Gen $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y\right)$;
$\sum_{i=1}^{N^{\prime}}-\log P\left(y_{i}\mid x,y_{<i}\right)$
$x_{i}$: image patch;
$y$: language instruction
Sim-Mani: Franka Kitchen;
Real-Mani (Franka): custom study desk
RPT [42] ViT [17, 41] MAE $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y,z,\dots\right)$ $x,y,z$: three distinct modalities Real-Mani (Franka): stack, pick, pick from bin
DINOv2 [49] ResNet, ViT Self-distillation $\sum_{x}\sum_{x^{\prime}\neq x}H(P_{t}(x),P_{x}(x^{\prime}))$ $x,x^{\prime}$: image views; $H()$: cross-entropy; $P_{t},P_{s}$: teacher, student (Used by OpenVLA [50], ReKep [51])
I-JEPA [52] ViT JEPA $\frac{1}{M}\sum_{i=1}^{M}\sum_{j}\left\lVert x_{j}(i)-y_{j}(i)\right\rVert_{2}^{2}$ $x_{j}(i)$: $j$-th masked patch embedding of block $i$; $y_{j}(i)$: unmasked patch embedding -
Theia [53] ViT-T/S/B Distillation (Distillation of vision foundation models, including ViT, CLIP, SAM, DINOv2, and Depth-Anything.) Sim-Mani&Navi: CortexBench [47]; Real-Mani: pick, place, open door/drawer
Voltron [48] introduces a pretraining objective by incorporating language conditioning and language generation into the MAE objective. Employing an encoder-decoder Transformer structure, the pretraining alternates between language-conditioned masked image reconstruction and language generation from masked images. This enhances the alignment between language and vision modalities.
VC-1 [47] conducts an in-depth examination of prior PVRs and introduces an enhanced PVR model by systematically exploring optimal ViT configurations across diverse datasets. Additionally, they perform a comprehensive comparative analysis of their model against previous methods on various manipulation and navigation datasets, shedding light on the critical factors that contribute to the improvement of PVR. Another study [54] also compares previous PVRs obtained under supervised learning or self-supervised learning.
DINOv2 [49] proposes a new self-supervised training paradigm for PVRs that achieves performance beyond that of MAE. It employs a self-distillation framework in which the teacher and student networks receive different views of the same image and match their encoded representations. The student network is updated using SGD, while the teacher network is maintained as the EMA of the student network.
I-JEPA [52] is motivated by the joint-embedding predictive architectures proposed by [55]. It constructs a “primitive” internal world model by comparing the embeddings of patches. Unlike DINO, which uses cropped images, I-JEPA employs masked patches. Moreover, it differs from MAE because it is a non-generative approach.
Theia [53] proposes distilling various vision foundation models into a single model. By fusing their information from segmentation, depth, semantics, etc., it outperforms previous PVRs while requiring less data and a smaller model size.
III-A3 Video Representations
Videos are simple sequences of images and can be represented by concatenating the usual PVRs of each frame. However, their multi-view nature enables a variety of unique representation techniques beyond those mentioned above, such as time contrastive learning and MAE. NeRF can be extracted from videos and contains rich 3D information for robot learning, as exemplified by F3RM [56] and 3D-LLM [57]. The recent 3D Gaussian Splatting (3D-GS) method has been gaining popularity in computer vision as well as in embodied AI. 3D Gaussians can be deformed through physics-based simulation for generative dynamics, as proposed in PhysGaussian [58], or they can serve as 3D representations for VLMs, as enabled by UniGS [59]. Additionally, many videos contain audio, which can provide important cues for robot policies [60].
III-A4 Dynamics Learning
Dynamics learning encompasses objectives aimed at endowing the model $f(\cdot)$ with an understanding of forward or inverse dynamics. Forward dynamics involves predicting the subsequent state resulting from a given action, whereas inverse dynamics entails determining the action required to transition from a previous state to a known subsequent state:
$\begin{split}\text{Forward dynamics: }&\hat{s}_{t+1}\leftarrow f_{\text{fwd}}(s_{t},a_{t}),\\ \text{Inverse dynamics: }&\hat{a}_{t}\leftarrow f_{\text{inv}}(s_{t},s_{t+1}).\end{split}$ (1)
Some approaches also frame these objectives as reordering problems for shuffled state sequences. Some forward dynamics methods closely resemble the image or video prediction pretraining used in PVRs. We compare them in Table II.
Table II: Dynamics learning methods for VLAs. $f(\cdot)$ is the dynamic model. Fwd, inv: forward & inverse dynamics.
Model V-Encoder Type Objective Notation Robotic Tasks
Vi-PRoM [61] ResNet Temporal dynamics $\sum_{i=1}^{T}\operatorname{CE}(i,f(x_{i}\mid x^{\prime}))$ $f(\cdot)$: predicts frame index in raw seq. $x$ given shuffled seq. $x^{\prime}$ Sim-Mani: Meta-World, Franka Kitchen; Real-Mani: open and close drawer/door
MaskDP [62] ViT-MAE [46] MAE (implicit fwd & inv dynamics) $\sum_{i=1}^{T}-\log P\left(x_{i}\mid x_{\neq i},y\right)$ $x_{i}$: state or action token; $y$: the other modality Sim: DeepMind Control Suite
MIDAS [63] ViT, Mask R-CNN [64] Inverse dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(a_{t},f_{\text{inv}}(s_{t},s_{t+1})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Mani: VIMA-Bench
SMART [65] CNN Forward & inverse dynamics $\sum_{i=1}^{T}\big{(}\operatorname{MSE}\left(s_{t+1},f_{\text{fwd}}(s_{t},a_{t})\right)$
$+\operatorname{MSE}\left(a_{t},f_{\text{inv}}(s_{t},s_{t+1})\right)\big{)}$
$s_{t}$, $a_{t}$: state, action Sim: DeepMind Control Suite
PACT [66] ResNet-18, PointNet [67] Forward dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(s_{t+1},f_{\text{fwd}}(s_{t},a_{t})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Navi: Habitat; Real-Navi (MuSHR vehicle)
VPT [68] ResNet Inverse dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(a_{t},f_{\text{inv}}(s_{t},s_{t+1})\right)$ $s_{t}$, $a_{t}$: state, action Sim: Minecraft
GR-1 [69] ViT-MAE Forward dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(s_{t+1},f_{\text{fwd}}(s_{t},a_{t})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Mani: CALVIN
Vi-PRoM [61] presents three distinct pretraining objectives. The first involves a contrastive self-supervised learning objective designed to distinguish between different videos. The remaining two objectives are centered around supervised learning tasks: temporal dynamics learning, aimed at recovering shuffled video frames, and image classification employing pseudo labels. Through a comprehensive comparison with preceding pretraining methods, Vi-PRoM demonstrates its effectiveness for behavior cloning and PPO.
MIDAS [63] introduces an inverse dynamics prediction task as part of its pretraining. The objective is to train the model to predict actions from observations, formulated as a motion-following task. This approach enhances the model’s understanding of the transition dynamics of the environment.
SMART [65] presents a pretraining scheme encompassing three distinct objectives: forward dynamics prediction, inverse dynamics prediction, and randomly masked hindsight control. The forward dynamics prediction task involves predicting the next latent state, while the inverse dynamics prediction task entails predicting the last action. In the case of hindsight control, the entire control sequence is provided as input, with some actions masked, and the model is trained to recover these masked actions. The incorporation of the first two dynamics prediction tasks facilitates capturing local and short-term dynamics, while the third task is designed to capture global and long-term temporal dependencies.
MaskDP [62] features the masked decision prediction task, wherein both state and action tokens are masked for reconstruction. This masked modeling task is specifically crafted to equip the model with an understanding of both forward and inverse dynamics. Notably, in contrast to preceding masked modeling approaches like BERT or MAE, MaskDP is applied directly to downstream tasks
PACT [66] introduces a pretraining objective aimed at modeling state-action transitions. It receives sequences of states and actions as input, and predicts each state and action token autoregressively. This pretrained model serves as a dynamics model, which can then be finetuned for various downstream tasks such as localization, mapping, and navigation.
VPT [68] proposes a video pretraining method that harnesses unlabeled internet data to pretrain a foundational model for the game of Minecraft. The approach begins by training an inverse dynamics model using a limited amount of labeled data, which is then utilized to label internet videos. Subsequently, this newly auto-labeled data is employed to train the VPT foundation model through behavior cloning. This methodology follows semi-supervised imitation learning. As a result of this process, the model demonstrates human-level performance across a multitude of tasks.
GR-1 [69] introduces video prediction pretraining tailored for a GPT-style model. The ability to anticipate future frames aligns with forward dynamics learning, contributing to more accurate action prediction.
III-A5 World Models
A world model $P(\cdot)$ encodes commonsense knowledge about the world and predicts the future state for a given action [55]:
$\begin{split}\hat{s}_{t+1}\sim P(\hat{s}_{t+1}|s_{t},a_{t}).\end{split}$ (2)
It enables model-based control and planning for embodied agents, as they can search for an optimal action sequence in imaginary space before executing any real actions. Although forward dynamics learning also attempts to predict the next state, it is usually treated as a pretraining task or an auxiliary loss to enhance the RL-Transformer-based action decoder for the primary robotics tasks, rather than serving as a standalone module. Additionally, new embodied experiences can be sampled from visual world models that explicitly generate images or videos of future states.
Dreamer [70] employs three primary modules to construct a latent dynamics model: a representation model, responsible for encoding images into latent states; a transition model, which captures transitions between latent states; and a reward model, predicting the reward associated with a given state. Under the actor-critic framework, Dreamer utilizes an action model and a value model to learn behavior through imagination by propagating analytic gradients through the learned dynamics. Building upon this foundation, DreamerV2 [71] introduces a discrete latent state space along with an improved objective. DreamerV3 [72] extends its focus to a broader range of domains with fixed hyperparameters. DayDreamer [73] applies this method to physical robots performing real-world tasks.
IRIS [74] employs a GPT-like autoregressive Transformer as the foundation of its world model, with a VQ-VAE serving as the vision encoder. Subsequently, a policy is trained using imagined trajectories, which are unrolled from a real observation by the world model. TWM [75] also investigates the application of Transformers in building world models.
III-A6 LLM-induced World Models
LLMs encompass a wealth of commonsense knowledge about the world, prompting many approaches to leverage that knowledge for improving VLAs.
DECKARD [76] prompts LLM to generate abstract world models (AWMs) represented as directed acyclic graphs, specifically tailored for the task of item crafting in Minecraft. DECKARD iterates between two phases: in the Dream phase, it samples a subgoal guided by the AWM; in the Wake phase, DECKARD executes the subgoal and updates the AWM through interactions with the game. This guided approach enables DECKARD to achieve markedly faster item crafting compared to baselines lacking such guidance.
LLM-DM [77] uses an LLM to construct world models in planning domain definition language (PDDL)—a capability that LLM+P [78] did not achieve, as its PDDL world model was hand-crafted. The LLM also acts as an interface, mediating between the generated PDDL model and corrective feedback from syntactic validators and human experts. Finally, the PDDL world model serves as a symbolic simulator, assisting the LLM planner in generating plans.
RAP [79] repurposes an LLM to act as both a policy that predicts actions and a world model that provides the state transition distribution. Unlike previous chain-of-thought prompting methods, RAP incorporates Monte Carlo Tree Search (MCTS) to enable structured planning, allowing the LLM to build a reasoning tree incrementally. This reasoning strategy helps RAP find a high-reward path that balances exploration and exploitation. Tree-Planner [80] improves efficiency by prompting the LLM only once to generate diverse paths within an action tree.
LLM-MCTS [81] builds upon RAP but extends the problem setting to POMDPs. As a world model, the LLM generates the initial belief of the current state; as a policy, it serves as a heuristic to guide action selection. By leveraging its commonsense knowledge, the LLM reduces the search space of MCTS, thereby improving search efficiency.
III-A7 Visual World Models
Unlike LLM-induced world models, which are in textual form, visual world models generate images, videos, or 3D scenes of future states—aligning more closely with the physical world. They can further be utilized to generate new trajectories. This direction has been gaining increasing attention since Sora demonstrated world-simulation capabilities [82], as investigated by a dedicated survey [83].
Genie [84] introduces a new class of generative models, termed Generative Interactive Environments. It consists of three main components: a spatiotemporal video tokenizer, an autoregressive dynamics model, and a latent action model. After being trained on unlabeled videos in an unsupervised manner, Genie allows users to interact with the generative environment on a frame-by-frame basis. Consequently, Genie establishes a foundation world model.
3D-VLA [85] proposes a 3D world model capable of goal generation. It processes visual inputs, such as images, depth map, and point clouds, and then generates a goal state—either as an image or a point cloud—using diffusion models in response to the user’s query. The generated goal state can subsequently be used to guide robot control.
UniSim [86] builds a generative model based on real-world interaction videos. It is capable of simulating visual outcomes for both high-level and low-level actions, which can then be leveraged as new experiences for training embodied agents. E2WM [87] even treats existing simulators as world models to collect embodied experiences via MCTS.
III-A8 Reasoning
Reasoning has become a key capability of LLMs, as demonstrated by chain-of-thought (CoT) methods [88, 89]. In embodied AI, researchers are exploring how to leverage CoT reasoning to refine the decision-making process.
ThinkBot [90] applies it to recover missing action descriptions in sparse human instructions. RAT [91] combines retrieval-augmented generation and CoT to elicit context-aware reasoning, improving long-horizon generation. ReAct [92] interleaves verbal reasoning traces and actions to allow for diverse types of reasoning. Tree-Planner [80] employs a tree-of-thoughts approach. ECoT [93] re-trains OpenVLA [50] using reasoning chain datasets, enabling embodied CoT reasoning that further enhances its performance.
Pros and Cons.
PVRs
Although time contrastive learning with videos and text-guided pretraining methods, such as CLIP, provide image-level information, they lack the pixel-level details offered by MAE-based self-supervised learning methods. Pixel-level information contains rich details—including segmentation masks, object positions, and depth estimation—that are generally more useful for robot manipulation tasks requiring high precision, as demonstrated by VC-1’s comparison [47]. DINOv2 learns both pixel- and image-level features by combining masked image modeling with a momentum encoder and multi-crop augmentation, with benefits on downstream robotic tasks evidenced by OpenVLA [50]. I-JEPA focuses on patches in the representation space and, as a result, captures low-level image features more effectively than view-invariance methods such as DINO. Theia distills various off-the-shelf vision foundation models into a single model that surpasses isolated individual models, as demonstrated by comprehensive evaluations against most existing PVRs in robot learning [53].
Forward & inverse dynamics
Forward dynamics learning is generally more challenging than inverse dynamics learning because predicting future states is more complex than predicting past actions. Consequently, the difficulty of forward dynamics often leads to greater performance improvements, as demonstrated in SMART. However, inverse dynamics models can be used to generate action labels for datasets that contain only states, such as raw robot manipulation videos.
World Models & Reasoning
Although both world models and reasoning methods can be applied to low-level control policies and high-level task planners, current approaches remain distinct. World models are predominantly used to interact with control policies because they excel at generating the immediate next state given low-level actions. In contrast, CoT-based reasoning methods focus on task planning since they express thought chains in text, making them well-suited for refining text-based task plans.
Figure 4: Illustration of a hierarchical robot policy. The high-level task planner decomposes the user instruction into subtasks, which are then executed step by step by the low-level control policy.
Table III: Low-level control policies. We also include some non-VLA models as they are closely related, marked in gray text. BC: behavior cloning (action type, cont/disc: continuous/discrete). TFM: Transformer. Xattn: cross-attention. Concat: concatenation. $p$/$s$: prompt/state vision encoder. [SC]: self-collect data. MPC: model predictive control.
Model V-Encoder L-Encoder VL-Fusion Action-Decoder Objective Training Dataset Robotic Tasks
CLIPort [43] CLIP-ResNet50, Transporter-ResNet CLIP-GPT Hadamard LingUNet BC ($\textbf{SE}(2)$) [SC] Sim: Ravens [94]
MCIL [95] Custom-CNN USE LMP RNN LMP RNN MCIL Play-LMP, [SC] Sim: 3D Playroom environment
HULC [96],
HULC++ [97]
MCIL CNN Sentence-BERT LMP TFM LMP RNN MCIL CALVIN data Sim: CALVIN
Language costs [98] CLIP-ViT,
UNet [99]
CLIP-GPT Concat STORM Max Likelihood Est. (cost map) [SC] Sim & Real (Franka): pick, place
Interactive Language [100] ResNet CLIP-GPT Xattn TFM BC (cont) Language-Table [SC] Sim & Real (xArm6): rearrangement
Hiveformer [101] UNet CLIP-GPT Concat TFM, UNet BC (cont) RLBench data Sim: RLBench [102]
PerAct [103] 3D CNN CLIP-GPT Concat PerceiverIO 3D affordance RLBench data Sim: RLBench; Real (Franka): put marker, place food, etc.
Act3D [104] CLIP-ResNet50 CLIP-GPT Xattn TFM 3D affordance (voxel) RLBench, (BC) Sim: RLBench; Real (Franka): reach target, wipe, stack, etc.
RVT, RVT-2 [105, 106] CLIP-ResNet50 CLIP-GPT Concat TFM 2D affordance (project to 3D) RLBench data, [SC] Sim: RLBench; Real (Franka): stack, press, place
RoboPoint [107] ViT-L/14 Vicuna-V1.5 13B [108] Concat - 2D affordance (project to 3D) [SC] Real (Franka): pick, place
Gato [36] ViT Sent.Piece Concat TFM BC (cont & disc) [SC] Sim & Real (Sawyer): RGB-stacking
RoboCat [109] VQ-GAN ($p,s$) [110] - - TFM BC, Observation prediction Self-improve Sim & Real (Sawyer, Franka, KUKA): stacking, building, lifting, insertion, removal
VIMA [111] ViT, Mask R-CNN [64] T5 [112] Xattn TFM BC ($\textbf{SE}(2)$) VIMA-Data [SC] Sim (Ravens): VIMA-Bench
BC-Z [113] ResNet18 ($p,s$) USE FiLM MLP BC (cont) [SC] Real (Everyday Robots): pick-place, pick-wipe, pick-drag, grasp, push
RT-1 [114] EfficientNet [115] USE FiLM TFM BC (disc) RT-1-Kitchen [SC] Real (Everyday Robots): pick-place, move, knock
MOO [111] OWL-ViT ($p$) [116], EfficientNet ($s$) USE FiLM TFM BC (disc) [SC] Real (Everyday Robots): pick, move near, knock, place upright, place into
Q-Transformer [117] EfficientNet USE FiLM TFM TD error Auto-collect + RT-1-Kitchen Sim: pick; Real (Everyday Robots): pick, place, open/close drawer, move near
RT-Trajectory [118] EfficientNet - - TFM BC (disc) [SC] Real (Everyday Robots): pick, place, fold towel, swivel chair, etc.
ACT [119] ResNet18 - - CVAE-TFM BC (cont, action chunking) [SC] with ALOHA Sim: transfer cube, bimanual insertion; Real (ViperX, WidowX): slot battery, open cup, etc
MT-ACT [120] (RoboAgent) CNN - FiLM CVAE-TFM BC (cont, action chunking) RoboSet Real (Franka): pick, place, open/close drawers, pour, push, drag, etc.
RoboFlamingo [121] CLIP-ViT-L/14 LLaMA [122], MPT [123] Xattn LSTM, TFM BC (cont) CALVIN data Sim: CALVIN
RoboUniView [124] UVFormer (Model design follows RoboFlamingo) BC (cont) CALVIN data Sim: CALVIN
VoxPoser [125] ViLD [126],
MDETR [127],
OWL-ViT [116],
SAM [18]
GPT-4 Code MPC - - Sim: Sapien; Real (Franka): move & avoid, set up table, close drawer, open bottle, sweep Trash
UniPi [128] Imagen video [129] T5-XXL Xattn CNN Inverse dynamics PF [SC], Ravens, BridgeV1 Sim: Painting Factory (PF), Ravens
Diffusion Policy [130] ResNet18 - - UNet,
TFM
DDPM [131] [SC] Sim: Robomimic, Franka Kitchen, etc; Real (UR5, Franka): push-T, flip mug, pour sauce
DP3 [132] DP3 Encoder (Model design follows Diffusion Policy) DDIM [SC] Sim: (72 tasks); Real (Franka, Allegro hand): roll-up, dumpling, drill, pour
SUDD [133] ResNet18 CLIP-GPT Concat UNet,
TFM
DDPM Lang-Guided Data Gen Sim: MuJoCo; Real (UR5e): pick, place
Octo [134] CNN T5-base Concat TFM DDPM OXE Real: BridgeV2, Stanford Coffee, Berkeley Peg Insert, etc.
3D Diffuser Actor [135] (Model design follows Act3D) DDPM RLBench, CALVIN, [SC] Sim: RLBench, CALVIN; Real (Franka): put, open, close, stack, etc.
MDT [136] CLIP-ViT-B/16 ($p$), Voltron/ResNet18 ($s$) CLIP-GPT Concat DiT [137] DDIM [138] CALVIN, LIBERO, [SC] Sim: CALVIN, LIBERO; Real (Franka): pick, push, open, close, etc.
RDT-1B [139] SigLIP T5-XXL Xattn DiT DDPM (Aggregated Datasets) Real (ALOHA dual-arm): wash, pour, fold, etc.
RT-2 [2] ViT-22B [140],
ViT-4B [141]
PaLI-X, PaLM-E Concat Symbol-tuning BC (disc), Co-fine-tuning VQA, RT-1-Kitchen Real: RT-1 evaluation tasks, Language-Table
RT-H [142] ViT-22B [140] PaLI-X 55B Concat Symbol-tuning BC (disc) Diverse+Kitchen Real: Diverse+Kitchen eval tasks
RT-X [143] (Models from RT-1 and RT-2) BC (disc) OXE [SC] Real: Kitchen Mani., NYU Door Open, BridgeV2, etc.
OpenVLA [50] DINOv2, SigLIP [144] Prismatic-7B [145] Concat Symbol-tuning BC (disc) OXE, DROID [146] Real: BridgeV2, RT-1 evaluation tasks
$\pi_{0}$ [147] SigLIP PaliGemma Concat Action expert Flow matching OXE, $\pi$-cross-embod. [SC] Real (general robot control): sweep, open, pack, etc.
III-B Low-level Control Policies
Through the integration of an action decoder with perception modules, such as vision encoders and language encoders, a VLA model $\pi_{\theta}$ with parameters $\theta$ is formed as a control policy to execute language instructions $p$:
$\begin{split}\hat{a}_{t}\sim\pi_{\theta}(\hat{a}_{t}|p,s_{\leq t},a_{<t}).\end{split}$ (3)
The diversity among control policy networks arises from individual modules and overall architectures. The most typical architecture is shown in Figure 1. This section explores various approaches to designing low-level control policies. Table III summarizes their technical details.
III-B1 Non-Transformer Control Policies
Prior to the adoption of Transformer models, early control policies for language-conditioned robotic tasks varied significantly in architecture.
CLIPort [43] integrates the vision and language encoders of CLIP [31] with the Transporter network [94], creating a two-stream architecture. In one stream, CLIP’s vision encoder extracts “semantic” information from the RGB image, while in the other stream, the Transporter network extracts “spatial” information from the RGB-D image. The CLIP sentence encoder encodes the language instruction and guides the output action, which is a pair of end-effector poses: the picking and placing poses. CLIPort demonstrates the ability to pick and place objects as specified by the language instruction.
BC-Z [113] processes two types of task instructions: a language instruction or a human demonstration video. The environment is presented to the model in the form of an RGB image. Then the instruction embedding and the image embedding are combined through the FiLM layer [148], culminating in the generation of actions. This conditional policy is asserted to exhibit zero-shot task generalization to unseen tasks.
MCIL [95] represents a pioneering robot policy that integrates free-form natural language conditioning. This is in contrast to earlier approaches that typically rely on conditions in the form of task id or goal image. MCIL introduces the capability to leverage unlabeled and unstructured demonstration data. This is achieved by training the policy to follow either image or language goals, with a small fraction of the training dataset consisting of paired image and language goals.
HULC [96] introduces several techniques aimed at enhancing robot learning architectures. These include a hierarchical decomposition of robot learning, a multimodal transformer, and discrete latent plans. The Transformer learns high-level behaviors, hierarchically dividing low-level local policies and the global plan. Additionally, HULC incorporates a visuo-lingual semantic alignment loss based on contrastive learning to align VL modalities. HULC++ [97] further integrates a self-supervised affordance model. This model guides HULC to the actionable region specified by a language instruction, enabling it to fulfill tasks within this designated area.
UniPi [128] treats the decision-making problem as a text-conditioned video generation problem. To predict actions, UniPi generates a video based on a given text instruction, and actions are extracted from the frames of the video through inverse dynamics. This innovative policy-as-video formulation offers several advantages, including enhanced generalization across diverse robot tasks and the potential for knowledge transfer from internet videos to real robots.
III-B2 Transformer-based Control Policies
Since the introduction of Transformers, control policies have converged to similar Transformer-based architectures.
Interactive Language [100] presents a robotic system wherein the low-level control policy can be guided in real-time by human instructions conveyed through language, enabling the completion of long-horizon rearrangement tasks. The efficacy of such language-based guidance is primarily attributed to the utilization of a meticulously collected dataset containing diverse language instructions, which surpasses previous datasets by an order of magnitude in scale.
Hiveformer [101] places significant emphasis on leveraging multi-view scene observations and maintaining the full observation history for a language-conditioned policy. This approach represents an advancement over previous systems such as CLIPort and BC-Z that only use the current observation. Notably, Hiveformer stands out as one of the early adopters of Transformer architecture as its policy backbone.
Gato [36] proposes a model that can play Atari games, caption images, and stack blocks, all with a single set of model parameters. This achievement is facilitated by a unified tokenization scheme, harmonizing the input and output across diverse tasks and domains. Consequently, Gato enables the simultaneous training of different tasks. Representing a significant milestone, Gato exemplifies the potential of constructing a “multi-modal, multi-task, multi-embodiment generalist agent”.
RoboCat [109] proposes a self-improvement process designed to enable an agent to rapidly adapt to new tasks with as few as 100 demonstrations. This self-improvement process iteratively finetunes the model and self-generates new data with the finetuned model. Built upon the Gato model, RoboCat incorporates the VQ-GAN image encoder [110]. During training, RoboCat predicts not only the next action but also future observations. The effectiveness of the self-improvement process is demonstrated through comprehensive experiments conducted in both simulated and real-world environments under multi-task, multi-embodiment settings.
RT-1 [114], developed by the same team as BC-Z, shares similarities with BC-Z but introduces some key distinctions. Notably, RT-1 employs a vision encoder based on the more efficient EfficientNet [115], departing from BC-Z’s use of ResNet. However, RT-1 does not use video as a task instruction. Additionally, RT-1 replaces the MLP action decoder in BC-Z with a Transformer decoder, producing discretized actions. This modification enables RT-1 to attend to past images, enhancing its performance compared to BC-Z.
Q-Transformer [117] extends RT-1 by introducing autoregressive Q-functions. In contrast to RT-1, which learns expert trajectories through imitation learning, Q-Transformer adopts Q-learning methods. Alongside the TD error objective of Q-learning, a conservative regularizer is incorporated to ensure that the maximum value action remains in-distribution. This approach allows Q-Transformer to leverage not only successful demonstrations but also failed trajectories for learning.
RT-Trajectory [118] adopts trajectory sketches as policy conditions instead of relying on language conditions or goal conditions. These trajectory sketches consist of curves that delineate the intended trajectory for the robot end-effector to follow. They can be manually specified through a graphical user interface, extracted from human demonstration videos, or generated by foundation models. RT-Trajectory’s policy is built upon the backbone of RT-1 and trained to control the robot arm to accurately follow the trajectory sketches. This approach facilitates generalization to novel objects, tasks, and skills, as trajectories from various tasks are transferable.
ACT [119] builds a conditional VAE policy with action chunking, requiring the policy to predict a sequence of actions rather than a single one. During inference, the action sequence is averaged using a method called temporal ensembling. RoboAgent [120] extends this approach with its MT-ACT model, demonstrating that action chunking improves temporal consistency. Additionally, RoboAgent introduces a semantic augmentation method that leverages inpainting to augment existing demonstrations.
RoboFlamingo [121] adapts the existing VLM, Flamingo [32, 149], to a robot policy by attaching an LSTM-based policy head to the VLM. This demonstrates that pretrained VLMs can be effectively transferred to language-conditioned robotic manipulation tasks.
III-B3 Control Policies for Multimodal Instructions
Multimodal instruction enables new ways to specify tasks, such as through demonstrations, by naming novel objects, or by pointing with a finger.
VIMA [150] places a significant emphasis on multimodal prompts and the generalization capabilities of models. By incorporating multimodal prompts, more specific and intricate tasks can be formulated compared to traditional pure text prompts. VIMA introduces four main types of tasks: object manipulation, visual goal reaching, novel concept grounding, one-shot video imitation, visual constraint satisfaction, visual reasoning. These tasks are often challenging or even infeasible to express using only language prompts. VIMA-Bench has been developed to evaluate across four generalizability levels: placement, combinatorial, novel object, novel task.
MOO [111] extends RT-1 to handle multimodal prompts. Leveraging the backbone of RT-1, MOO incorporates OWL-ViT to encode images within the prompt. By expanding the RT-1 dataset with new objects and additional prompt images, MOO enhances the generalization capabilities of RT-1. This extension also facilitates new methods of specifying target objects, such as pointing with a finger, and clicking on graphical user interfaces.
III-B4 Control Policies with 3D Vision
We live in a 3D world, and it is intuitive that using 3D representations as visual inputs should provide richer information than 2D images. Point clouds are a popular choice for representing 3D inputs due to their straightforward derivation from RGBD inputs, as demonstrated by both DP3 and 3D Diffuser Actor. However, voxels have also been explored in various works. RoboUniView [124] shows improved performance by injecting 3D information into RoboFlamingo through a novel UVFormer module as its vision encoder, which provides 3D occupancy information from multi-perspective images. VER [151] also proposes voxelizing multi-view images into 3D cells in a coarse-to-fine manner, enhancing performance in vision-language navigation tasks.
PerAct [103] represents an advancement in both the observation and action spaces by leveraging 3D voxel representations. This approach offers a robust structural prior for action learning, enabling natural processing of multi-view observations and facilitating data augmentation in 6-DoF. In this framework, the input to the model comprises voxel maps reconstructed from RGBD images, while the output corresponds to the best voxel for guiding the gripper’s movement. By adopting this formulation, PerAct facilitates efficient task learning even with a few demonstrations. On the other hand, Act3D [104] introduces a continuous resolution 3D feature field, with adaptive resolutions based on current task, addressing the computational cost of voxelization. RVT, RVT-2 [105, 106] propose re-rendering images from virtual views of the scene point cloud and using these images as inputs, rather than directly relying on 3D inputs.
III-B5 Diffusion-based Control Policies
Diffusion-based action generation leverages the success of diffusion models in the field of computer vision.
Diffusion Policy [130] formulates a robot policy as a DDPM [131]. This approach incorporates a variety of techniques, including receding horizon control, visual conditioning, and the time-series diffusion transformer. The effectiveness of this diffusion-based visuomotor policy is underscored by its proficiency in multimodal action distributions, high-dimensional action spaces, and training stability.
SUDD [133] presents a framework where an LLM guides data generation, and subsequently, the filtered dataset is distilled into a visuo-linguo-motor policy. This framework achieves language-guided data generation by composing an LLM with a suite of primitive robot utilities, such as grasp samplers and motion planners. It then extends Diffusion Policy by incorporating language-based conditioning for multi-task learning and facilitates the distillation of the filtered dataset.
Octo [134] introduces a transformer-based diffusion policy characterized by a modular open-framework design, allowing for flexible connections from different task definition encoders, observation encoders, and action decoders to the Octo Transformer. Being among the first to utilize the Open X-Embodiment dataset [143], Octo demonstrates positive transfer and generalizability across diverse robots and tasks.
MDT [136] adapts the newly introduced DiT model [137] from computer vision to the action prediction head. DiT, originally proposed as a transformer-based diffusion model, replaces the classical U-Net architecture for video generation. Coupled with two auxiliary objectives—masked generative foresight and contrastive latent alignment—MDT demonstrates better performance than the U-Net-based diffusion model, SUDD.
RDT-1B [139] is a diffusion-based foundation model for bimanual manipulation, also built on DiT. It addresses data scarcity by introducing a unified action format across various robots, enabling pretraining on heterogeneous multi-robot datasets with over 6K trajectories. As a result, RDT scales up to 1.2B parameters and demonstrates zero-shot generalization.
III-B6 Diffusion-based Control Policies with 3D Vision
Several works have proposed combining 3D Vision with diffusion-based policies. DP3 [132] introduces 3D inputs to a diffusion policy, resulting in improved performance. Similarly, 3D Diffuser Actor [135] shares the core idea of DP3 but differs in model architecture by combining Act3D with Diffusion Policy.
III-B7 Control Policies for Motion Planning
Motion planning involves decomposing movement tasks into discrete waypoints while satisfying constraints such as obstacle avoidance and kinematic limits.
Language costs [98] presents a novel approach to robot correction using natural language for human-in-the-loop robot control systems. This method leverages predicted cost maps generated from human instructions, which are then utilized by the motion planner to compute the optimal action. This framework enables users to correct goals, specify preferences, or recover from errors through intuitive language commands.
VoxPoser [125] employs LLM and VLM to create two 3D voxel maps that represent affordance and constraint. It leverages the programming capability of LLMs, and the perception capability of VLMs. LLM translates language instructions into executable code, invoking VLM to obtain object coordinates. Based on the composed affordance and constraint maps, VoxPoser employs model predictive control to generate a feasible trajectory for the robot arm’s end-effector. Notably, VoxPoser does not require any training, as it directly connects LLM and VLM for motion planning.
RoboTAP [152] breaks down demonstrations into stages marked by the opening and closing of the gripper. In each stage, RoboTAP uses the TAPIR algorithm to detect active points that track the relevant object from the source to the target pose. The path can then be used by visual servoing to control the robot. A motion plan is created by stringing together these stages, enabling few-shot visual imitation.
III-B8 Control Policies with Point-based Action
Recent research has explored leveraging the capabilities of VLMs to select or predict point-based actions—a cost-effective alternative to building full VLAs.
PIVOT [153] casts robotics tasks as visual question answering, leveraging VLMs to select the best robot action from a set of visual proposals. Visual proposals are annotated in the form of keypoints on images. The VLM is iteratively prompted to refine them until the best option is identified.
RoboPoint [107] finetunes VLM using the task of spatial affordance prediction, which is to point at where to act on the image. These affordance points on 2D images are subsequently projected into 3D space using depth maps, forming the predicted robot action.
ReKep [51] is a constraint function that maps 3D keypoints in the scene to a numerical cost. Robotics manipulation tasks can be represented as a sequence of ReKep constraints, which are produced with large vision models and VLMs. Consequently, robot actions can be obtained by solving the constrained optimization problem.
III-B9 Large VLA
Large VLA is equivalent to the original VLA definition proposed by RT-2 [2], as illustrated in Figure 2a. This terminology is analogous to the distinction between LLMs and general language models, or between large VLMs and general VLMs.
RT-2 [2] endeavors to harness the capabilities of large multimodal models in robotics tasks, drawing inspiration from models like PaLI-X and PaLM-E. The approach introduces co-fine-tuning, aiming to fit the model to both Internet-scale visual question answering (VQA) data and robot data. This training scheme enhances the model’s generalizability and brings about emergent capabilities.
RT-H [142] introduces an action hierarchy that includes an intermediate prediction layer of language motions, situated between language instructions and low-level actions (translation and rotation). This additional layer facilitates improved data sharing across different tasks. For example, both the language instructions “pick” and “pour” may involve the language motion “move the arm up”. Moreover, this action hierarchy enables users to specify corrections to recover from failures, which the model can then learn from.
RT-X [143] builds upon the previous RT-1 and RT-2 models. These models are re-trained using the newly introduced open-source large dataset, named Open X-Embodiment (OXE), which is orders of magnitude larger than previous datasets. The resulting models, RT-1-X and RT-2-X, both outperform their original versions.
OpenVLA [50] was later developed as an open-source counterpart to RT-2-X. They additionally explored efficient fine-tuning methods, including LoRA and model quantization.
$\pi_{0}$ [147] proposes a flow-matching architecture for transforming VLMs into VLAs. By incorporating an additional action expert based on the mixture-of-experts framework, it effectively inherits the Internet-scale knowledge in the VLM while extending its capabilities to address robotic tasks.
(a) FiLM
(b) Cross-attention
(c) Concatenation
Figure 5: Three common vision-language fusion methods of Transformer-based control policies. FiLM layers (Hadamard product $\odot$) are used in RT-1 models. Some utilize cross-attention to condition on instructions. Concatenation ($\oplus$) is the prevailing method in LVLAs.
Pros and Cons.
Architectures
Various VLA architectures explore different methods of fusing vision and language inputs, including cross-attention, FiLM, and concatenation, as illustrated in Figure 5. FiLM is used in RT-1 and thus its follow-up works inherit this mechanism. While cross-attention may offer superior performance with smaller model sizes, concatenation is simpler to implement and can achieve comparable results with larger models [150].
Action types and their training objectives
Most low-level control policies predict actions for the end-effector pose while abstracting away the motion planning module that produces more fine-grained motions. While this abstraction facilitates better generalization to different embodiments, it also imposes limitations on dexterity.
The behavior cloning (BC) objective is used in imitation-learning, with different variants for different action types. The BC objective for continuous action [121] can be written as:
$\mathcal{L}_{\text{Cont}}=\sum_{t}\operatorname{MSE}\left(a_{t},\hat{a}_{t}\right),$ (4)
where $\operatorname{MSE}(\cdot)$ stands for mean squared error. $a_{t}$ is the action annotation from expert demonstrations.
Discrete action is achieved by dividing the action value range into a number of bins [114]. Its BC objective is:
$\mathcal{L}_{\text{Disc}}=\sum_{t}\operatorname{CE}\left(a_{t},\hat{a}_{t}\right),$ (5)
where $\operatorname{CE}(\cdot)$ stands for cross-entropy loss.
CLIPort [43] and VIMA [150] use $SE(2)$ action and its BC objective can be expressed as follows:
$\begin{split}\mathcal{L}_{SE(2)}&=\operatorname{CE}(a_{\text{pick}},\hat{a}_{\text{pick}})+\operatorname{CE}(a_{\text{place}},\hat{a}_{\text{place}})\end{split}$ (6)
The DDPM objective in diffusion-based control policies [130] is represented as:
$\mathcal{L}_{\text{DDPM}}=\operatorname{MSE}\left(\varepsilon^{k},\varepsilon_{\theta}(a_{t}+\varepsilon^{k},k)\right)$ (7)
where $\varepsilon^{k}$ is a random noise for iteration $k$; $\varepsilon_{\theta}$ is the noise prediction network, i.e., the VLA model.
While discrete action has demonstrated superior performance in RT-1 [114], Octo [134] argues that it leads to early grasping issues. $SE(2)$ actions require the model to predict only the pick and place end-effector poses, which is sufficient for many tabletop manipulation tasks. However, more complex tasks—such as “pouring water into a cup”—may require additional DoFs, thereby necessitating $SE(3)$ actions [51]. Although point-based actions can be coarse-grained, they are more easily obtained from VLMs in a zero-shot manner [51].
Table IV: High-level task planners. VL: vision-language fusion method.
Model V-Encoder Lang Model VL Control Policy Train Robotic Tasks
SayCan [10] - PaLM, FLAN - BC-Z, MT-Opt No Real-Navi&Mani (Everyday): office kitchen
[…]
(SL)3 [155] - T5-small - seq2seq Yes Sim-Navi: ALFRED
EmbodiedGPT [156] EVA ViT-G/14 [157] LLaMA-7B Embodied-former MLP Yes Sim-Mani: Franka Kitchen, Meta-World
PaLM-E [11] ViT, OSRT [158] PaLM Concat Interactive Language, RT-1 Yes Sim: TAMP; Sim&Real-Mani: Lang.-Table; Real: SayCan setup
LEO [159] ConvNeXt [160], PointNet++ [161] Vicuna-7B Concat MLP Yes Sim-Mani: CLIPort task subset; Sim-Navi: Habitat-web
3D-LLM [57] CLIP, 3D-CLR [162], ConceptFusion [163] Flamingo, BLIP-2 Concat DD-PPO [164] Yes Sim-Navi: Habitat
ShapeLLM [165] ReCon++ LLaMA Concat - Yes 3D MM-Vet Benchmark
Inner Monologue [9] MDETR, CLIP InstructGPT, PaLM Language CLIPort, MDETR, SayCan policies No Sim&Real-Mani: Tabletop Rearrangement; Real-Navi&Mani: Kitchen Mobile Mani.
LLM-Planner [166] HLSM GPT-3 Language HLSM Yes Sim-Navi: ALFRED
LID [167] CNN GPT-2 Language & Embed - Yes Sim: VirtualHome
SMs [168] CLIP, ViLD GPT-3, RoBERTa [169] Language & Code CLIPort No Sim-Mani: pick, place
ConceptGraphs [170] SAM, CLIP, LLaVA [8] GPT-4 Code (JSON) API No Sim: AI2-THOR; Real (Spot Arm): pick, place
ProgPrompt [171] ViLD GPT-3 Code API No Sim: VirtualHome
ChatGPT for Robotics [172] YOLOv8 [173] ChatGPT Code API No Real-Navi: drone flight; Sim-Navi: AirSim, Habitat
CaP [174] ViLD, MDETR GPT-3, Codex Code API No Sim-Mani & Sim-Navi: pick, place, etc.
DEPS [175] CLIP ChatGPT Code MC-Controller, Steve-1 No Sim: Minecraft
RT series
RT-1 [114] inspired a series of “Robotic Transformer” models. The Transformer backbone surpasses previous RNN backbones by harnessing the higher capacity of Transformers to absorb larger robot datasets. Preceding RT-1 was BC-Z, which solely utilized MLP layers for action prediction. Subsequent to RT-1, several works emerged, each introducing new capabilities. MOO adapted RT-1 to accommodate multimodal prompts. RT-Trajectory enabled RT-1 to process trajectory sketches as prompts. Q-Transformer utilized Q-learning to train RT-1. RT-2, based on ViT and LLM, introduced a completely different architecture from RT-1. RT-X retrained RT-1 and RT-2 with a significantly larger dataset, resulting in improved performance. Based on RT-2, RT-H [142] introduces action hierarchies for better data sharing.
LVLA vs VLA
While LVLAs can greatly enhance instruction-following abilities because they can better parse user intentions, concerns arise regarding their training cost and deployment speed. Slow inference speed, in particular, can significantly impact performance in dynamic environments, as changes in the environment may occur during inference. Therefore, several methods have been proposed to improve efficiency. TinyVLA [176] focused on inference speed and data efficiency through smaller VLM and diffusion head for robot action. DeeR-VLA [177] proposes only partially activating the model with dynamic inference with early-exit.
IV Task Planners
A high-level task planner $\pi_{\phi}$ aims to divide a complex task $\ell$ into a sequence of subtasks $p_{i}$ (i.e., task plan), each serving as a language instruction to low-level control policies $\pi_{\theta}$:
$\begin{split}\left[p_{1},p_{2},\dots,p_{N}\right]\sim\pi_{\phi}(\ell,s_{t}).\end{split}$ (8)
This process is often referred to as task decomposition and is also related to the task and motion planning (TAMP) setting. Ideally, task plans should also incorporate optimal scheduling for these subtasks. When equipped with task planners, VLAs can complete more complex, long-horizon tasks, as illustrated in Figure 4. Table IV outlines important details about task planners.
IV-A Monolithic Task Planners
A single LLM or multimodal LLM (MLLM) can typically generate task plans by employing a tailored framework or through finetuning on embodied datasets. We refer to these as monolithic models.
IV-A1 Grounded Task Planners
Grounded task planning involves generating high-level actions while considering whether they can be executed by low-level control policies.
SayCan [10] is a framework designed to integrate high-level LLM planners with low-level control policies. In this framework, the LLM planner accepts users’ high-level instruction and “says” what the most probable next low-level skill is, a concept referred to as task-grounding. The low-level policy provides the value function as the affordance function, determining the possibility that the policy “can” complete the skill, known as world-grounding. By considering both LLM’s plan and affordance, the framework selects the optimal skill for the current state.
[…]
(SL)3 [155] is a learning algorithm that alternates between three steps: segmentation, labeling, and parameter update. In the segmentation step, high-level subtasks are aligned with low-level actions, subtask descriptions are then inferred in the labeling step, and finally, the network parameters are updated. This approach enables a hierarchical policy to discover reusable skills with sparse natural language annotations.
IV-A2 End-to-end Task Planners
Similar to LVLAs, task planners can be implemented as end-to-end multimodal LLMs, leveraging their Internet-scale knowledge for task planning. Clearly, they are also monolithic models.
PaLM-E [11] integrates ViT and PaLM to create a large embodied multimodal language model capable of performing high-level embodied reasoning tasks. Based on perceived images and high-level language instructions, PaLM-E generates a text plan that serves as instructions for low-level robotic policies. In a mobile manipulation environment, they map the generated plan to executable low-level instructions with SayCan [10]. As the low-level policy executes actions, PaLM-E can also replan based on changes in the environment. With PaLM as its backbone, PaLM-E can handle both normal visual question answering (VQA) tasks, along with the additional embodied VQA tasks.
EmbodiedGPT [156] introduces the embodied-former, which outputs task-relevant instance-level features. This is achieved by incorporating information from vision encoder embeddings and embodied planning information provided by LLM. The instance feature serves to inform the low-level policy about the immediate next action to take.
IV-A3 End-to-end Task Planners with 3D Vision
Some task planners also explore the use of 3D vision. Because the majority of current MLLMs deal with images as visual inputs, they require architectural changes to incorporate 3D visual inputs, and therefore they are usually end-to-end models.
LEO [159] uses a two-stage training strategy to integrate a point cloud encoder with an LLM: the first stage focuses on 3D vision-language alignment, followed by the second stage, which involves 3D vision-language-action instruction tuning. LEO demonstrates proficiency not only in 3D question-answering tasks but also in manipulation, navigation, and task planning.
3D-LLM [57] injects 3D information into LLMs and empowers them to perform 3D tasks, such as 3D-assisted dialog, and navigation. The 3D features can come in diverse forms, including point cloud, gradSLAM, and neural voxel field. MultiPLY [179] is an object-centric embodied LLM that incorporates even more modalities, including audio, tactile, and thermal.
ShapeLLM [165] is built on the novel 3D vision encoder ReCon++, which distills knowledge from multi-view image and text teachers, along with a point cloud MAE. By integrating ReCon++ with LLaMA, ShapeLLM improves embodied interaction performance on their newly proposed 3D benchmark, 3D MM-Vet.
(a) Language-based
(b) Code-based
Figure 6: Different approaches to connect LLM to multi-modal modules in modular task planners.
IV-B Modular Task Planners
Finetuning end-to-end models on embodied data can be expensive, and there are approaches that adopt a modular design by assembling off-the-shelf LLMs and VLMs into task planners.
IV-B1 Language-based Task Planners
Language-based approaches use natural language descriptions as the medium for exchanging multimodal information, as illustrated in Figure 6a.
Inner Monologue [9] sits between high-level command and low-level policy to enable closed-loop control planning. It employs LLM to generate language instructions for low-level control policies and dynamically updates these instructions based on feedback received from the control policies. The feedback encompasses various sources: success feedback, object and scene feedback, and human feedback. As the feedback is communicated to the LLM in textual format, no additional training is required for the LLM. A similar approach is used in ReAct [92].
LLM-Planner [166] introduces a novel approach to constructing a hierarchical policy comprising a high-level planner and a low-level planner. The high-level planner harnesses the capabilities of LLM to generate natural language plans, while the low-level planner translates each subgoal within the plan into primitive actions. While sharing similarities with previous methods in its overall architecture, LLM-Planner distinguishes itself by incorporating a re-planning mechanism, aiding the robot to “get unstuck”.
LID [167] introduces a novel data collection procedure termed Active Data Gathering (ADG). A key aspect of ADG is hindsight relabeling, which reassigns labels to unsuccessful trajectories, effectively maximizing the utilization of data irrespective of their success. By converting all environmental inputs into textual descriptions, their language-model-based policy demonstrates enhanced combinatorial generalization.
Socratic Models (SMs) [168] presents a unique framework wherein diverse pretrained models are effectively composed without the need for finetuning. The framework is based on the key component, named multimodal-informed prompting, facilitating information exchange among models with varied multimodal capabilities. The idea is to utilize multimodal models to convert non-language inputs into language descriptions, effectively unifying different modalities within the language space. Beyond excelling in conventional multimodal tasks, SMs showcases its versatility in robot perception and planning. In addition to natural language plans, task plans can also be represented in the form of pseudocode.
Table V: Recent datasets for robot learning. Following the definition in RT-X, skills correspond to verbs, and tasks are different combinations of verbs and objects. $\ast$ Dataset collected in simulators rather than in the real world. Some datasets are continuously updated; this table is based on the information from their original papers. Table adapted from [180, 146].
Name Collection Visual Instruction Embodiment Skills Tasks Scenes Objects Episodes
MIME [181] Human Teleoperation RGB, D Demo Baxter 12 20 1 - 8.3K
RoboTurk [182] $\ast$ Human Teleoperation RGB - Sawyer 2 - 1 - 2.1K
RoboNet [183] Scripted RGB Goal image (7 robots) - - 10 - 162K
MT-Opt [184] Scripted RGB Lang (7 robots) 2 12 1 - 800K
BC-Z [113] Human Teleoperation RGB Lang, Demo Everyday 3 100 1 - 25.9K
RT-1-Kitchen [114] Human Teleoperation RGB Lang Everyday 12 700+ 2 16 130K
MOO [111] Human Teleoperation RGB Multi-modal Everyday 5 - - 106 59.1K
VIMA [150] $\ast$ Scripted RGB Multi-modal UR5 17 - 1 29 650K
RoboSet [185] Human, Scripted RGB, D Lang Franka Panda 12 38 11 - 98.5K
BridgeData V2 [186] Human, Scripted RGB, D Lang WidowX 250 13 - 24 100+ 60.1K
RH20T [180] Human Teleoperation RGB, D Lang (4 robots) 42 147 7 - 110K+
DROID [146] Human Teleoperation RGB, D Lang Franka Panda 86 - 564 - 76K
OXE [143] Aggregated Datasets RGB, D Lang (22 robots) 527 160,266 311 - 1M+
IV-B2 Code-based Task Planners
Code-based task planners leverage the coding ability of LLMs to generate task plans in the form of a program. Object detectors, VLMs, and control policies can be invoked via APIs, as shown in Figure 6b.
ProgPrompt [171] introduces a novel task-planning approach by prompting LLMs with program-like specifications detailing available actions and objects. This enables LLMs to generate high-level plans for household tasks in a few-shot manner. Environmental feedback can be incorporated through assertions within the program.
ChatGPT for Robotics [172] take advantage of the programming ability of ChatGPT to facilitate “user on the loop” control, a departure from the conventional “engineer in the loop” methodology. The procedure includes several steps: firstly, a list of APIs is defined, such as an object-detection API, a grasp API, a move API; secondly, a prompt is then constructed for ChatGPT, specifying the environment, API functionality, task goal, etc.; thirdly, iteratively prompting ChatGPT to write code with the defined APIs that can execute the task, provided the access to simulation and user feedback for evaluating the code quality and safety; finally, executing the ChatGPT generated code. In this procedure, ChatGPT serves as a high-level task planner and actions are generated through function calls to corresponding low-level APIs.
Code as policies (CaP) [174] also leverages the code-writing capability of LLMs. It employs GPT-3 or Codex to generate policy code, which, in turn, invokes perception modules and control APIs. CaP exhibits proficiency on spatial-geometric reasoning, generalization to new instructions, and parameterization for low-level control primitives. By leveraging the multimodal capabilities of GPT-4V, COME-robot [187] eliminates the need for perception APIs in CaP. This also opens up possibilities for open-ended reasoning and adaptive planning within a closed-loop framework, enabling capabilities such as failure recovery and free-form instruction following.
DEPS [175] stands for “Describe, Explain, Plan and Select”. This approach employs an LLM to generate plans and explain failures based on feedback descriptions collected from the environment—a process referred to as “self-explanation”, aiding in re-planning. Additionally, DEPS introduces a trainable goal selector to choose among parallel candidate sub-goals based on how easily they can be achieved, a crucial aspect often overlooked by other high-level task planners.
ConceptGraphs [170] introduces a method to convert observation sequences into open-vocabulary 3D scene graphs. Objects are extracted from RGB images using 2D segmentation models, and VLMs are employed to caption objects and establish inter-object relations, resulting in the formation of the 3D scene graph. This graph can then be translated into a text description (JSON), offering rich semantic and spatial relationships between entities to LLMs for task planning.
Pros and Cons. Monolithic task planners that utilize grounded task planning focus on generating executable plans. End-to-end models like PaLM-E share an architecture similar to most LVLAs and can be finetuned on specialized embodied data to achieve better performance. However, the training costs of such large models can be a significant concern. In contrast, modular task planners are more readily deployable because they leverage off-the-shelf LLMs and VLMs. Language-based task planners offer the advantage of seamless integration of LLMs and VLMs, as they are designed to operate in the natural language space. However, they often require extra steps to align the generated task plans with language instructions that are admissible to low-level control policies. On the other hand, while code-based task planners may require manually wrapping VLMs and control policies in APIs and preparing clear documentation in advance, they enable code debugging and provide greater controllability. Nevertheless, their performance can be constrained by the programming capabilities of existing models.
V Datasets and Benchmarks
Table VI: Simulators and simulated benchmarks. Control: Continuous control tasks. D, S, A, N: Depth, segmentation, audio, normals. Force: simulated contact force between end-effector and item. PD: Pre-defined. Table adapted from [188, 189].
Name Scenes
/Rooms
Objects
/Cat
UI Physics Engine Task Observation Action Agent Description Related
Gibson [190] 572/- - - Pybullet Navi RGB, D, N, S - - Navi only -
iGibson [191, 192, 189] 15/108 152/5 Mouse, VR Pybullet Navi, Mani RGB, D, S, N, Flow, LiDAR Force TurtleBot v2, LoCoBot, etc. VR, Continuous Extended States. Versions: iGibson 0.5, 1.0, 2.0 Benchmarks: BEHAVIOR-100 [188], BEHAVIOR-1K [193]
SAPIEN [194] - 2346/- Code PhysX Navi, Mani RGB, D, S Force Franka Articulation, Ray Tracing VoxPoser. Benchmarks: SIMPLER [195]
AI2-THOR [196] -/120 118/118 Mouse Unity Navi, Mani RGB, D, S, A Force, PD ManipulaTHOR, LoCoBot, etc. Object States, Task Planning. Versions: [197, 198] Benchmarks: ALFRED, RoPOR [199]
VirtualHome [200] 7/- -/509 Lang Unity Navi, Mani RGB, D, S Force, PD Human Object States, Task Planning LID, Translated〈LM〉, ProgPrompt
TDW [201] 15/120 112/50 VR Unity, Flex Navi, Mani RGB, D, S, A Force Fetch, Sawyer, Baxter Audio, Fluids -
RLBench [102] 1/- 28/28 Code Bullet Mani RGB, D, S Force Franka Tiered Task Difficulty Hiveformer, PerAct
Meta-World [202] 1/- 80/7 Code MuJoCo Mani Pose Force Sawyer Meta-RL R3M, VC-1, Vi-PRoM, EmbodiedGPT
CALVIN [203] 4/- 7/5 - Pybullet Mani RGB, D Force Franka Long-horizon Lang-cond tasks GR-1, HULC, RoboFlamingo
Franka Kitchen [204] 1/- 10/6 VR MuJoCo Mani Pose Force Franka Extended by R3M with RGB R3M, Voltron, Vi-PRoM, Diffusion Policy, EmbodiedGPT
Habitat [205, 206] Matterport + Gibson Mouse Bullet Navi RGB, D, S, A Force Fetch, Franka, AlienGO Fast, Navi only. Versions: Rearrangement [207], Habitat 2.0 VC-1, PACT; OVMM [208]
ALFRED [209] -/120 84/84 - Unity Navi, Mani RGB, D, S PD Human Diverse long-horizon tasks (SL)3, LLM-Planner
DMC [210] 1/- 4/4 Code MuJoCo Control RGB, D Force - Continuous RL VC-1, SMART
OpenAI Gym [211] 1/- 4/4 Code MuJoCo Control RGB Force - Single agent RL environments -
Genesis [212] (Rigid, deformable, liquid, etc.) Code (Propri-etary) Navi, Mani RGB, D, S, N Force Franka, Unitree, etc. High-speed comprehensive physics simulation -
V-A Real-world Robot Datasets & Benchmarks
Embodied AI faces significant data scarcity issues because real-world robot data is not as readily available as language data. Collecting real-world robot datasets poses multiple challenges. Firstly, it is impeded by the cost and time required to procure robotic equipment, set up environments, and gather expert data through dedicated policies or human teleoperation. Secondly, the diverse types and configurations of robots introduce inconsistencies in sensory data, control modes, gripper types, etc. Lastly, accurately capturing object 6D poses and reproducing setups remains elusive. We summarize recent robot datasets in Table V. In addition, real-world benchmarks are further complicated by the need for human evaluation.
V-B Simulators, Simulated Robot Datasets & Benchmarks
Many researchers resort to simulated environments to circumvent real-world obstacles and scale the data collection process. We compare simulators and simulated benchmarks in Table VI. Nevertheless, this strategy presents its own challenges, chief among which is the sim-to-real gap. This discrepancy arises when models trained on simulated data exhibit poor performance during real-world deployment. The causes of this gap are multifaceted, encompassing unrealistic rendering quality, inaccuracies in physics simulations, and domain shifts characterized by object properties and robot motion planners. For instance, simulating non-rigid objects such as deformable objects or liquids presents significant difficulties. Moreover, importing new objects into simulators requires considerable effort, often involving techniques such as 3D scanning and mesh editing. Despite these hurdles, simulated environments provide automated evaluation metrics that aid researchers in consistently evaluating robotic models. Many benchmarks are based on simulators because they can precisely reproduce the experimental setup and yield fair comparisons of different models. Another technique, known as real-to-sim, can improve simulation fidelity, recreate failure cases, or facilitate digital twins.
V-C Automated Dataset Collection
Several approaches advocate for automated dataset collection. RoboGen [213] employs a generative simulation paradigm that proposes interesting skills, simulates corresponding environments, and selects optimal learning approaches to train policies for acquiring those skills. AutoRT [214] functions as a robot orchestrator driven by LLMs, generating tasks, filtering them by affordance, and utilizing either autonomous policies or human teleoperators to collect and evaluate data. DIAL [215] focuses on augmenting language instructions in existing datasets using VLMs. RoboPoint [107] generates scenes procedurally with randomized 3D layouts, objects, and camera viewpoints.
V-D Human Datasets
An alternative strategy to address data scarcity in real-world settings is to leverage human data. Human behavior offers plentiful guidance for robot policies due to its dexterity and diversity. However, this strategy also comes with inherent drawbacks. Capturing and transferring human hand/body motions to robot embodiments is inherently difficult. Moreover, the inconsistency in human data poses a hurdle, as some data may be egocentric while others are captured from third-person perspectives. Additionally, filtering human data to extract useful information can be labor-intensive. These obstacles underscore the complexities involved in incorporating human data into robot learning processes. UMI [216] proposes a method to mitigate these issues using hand-held grippers. For a more comprehensive comparison of human datasets, we refer interested readers to [217].
V-E Task Planning Benchmarks
EgoPlan-Bench [218] focuses on benchmarking real-world task planning with human annotations. PlanBench [219, 220] comprehensively assesses various aspects of task planning ability, such as cost optimality, plan verification, and replanning. LoTa-Bench [221] directly evaluates task planning by executing the generated plans in simulators and calculating success rates. Embodied Agent Interface (EAI) [222] argues that this approach fails to pinpoint issues in LLMs. By formalizing the input-output of LLM-based modules for decision-making tasks, EAI enables more fine-grained metrics beyond success rates.
V-F Embodied Question Answering Benchmarks
Embodied question answering (EQA) benchmarks, as summarized in Table VII, do not directly evaluate robotic tasks like manipulation and navigation, but they are aimed at other relevant abilities for embodied AI, such as spatial reasoning, physics understanding, as well as world knowledge. EQA is akin to previous visual question answering benchmarks, but differs in that the agent can actively explore the environment before providing an answer. EmbodiedQA [223] and IQUAD [224] were among the first works to introduce this type of benchmarks. MT-EQA [225] focuses on complex questions involving multiple targets. MP3D-EQA [226] converts previous RGB inputs to point clouds, testing 3D perception capabilities.
Active exploration requires access to a simulator, limiting the types of data that can be used, such as real-world videos. EgoVQA [227] shifts the focus of VQA to egocentric videos. EgoTaskQA [228] emphasizes spatial, temporal, and causal relationship reasoning. EQA-MX [229] investigates multimodal expressions (MX), including regular verbal utterances and nonverbal gestures like eye gaze and pointing. OpenEQA [230] evaluates seven main categories, including functional reasoning and world knowledge.
Table VII: Embodied question answering benchmarks. Acc: accuracy. PPL: perplexity.
# QA # Videos Video source Answer type Active Data collection Metrics
EQA [223] 5K 750 envs House3D simulator Answer set 172 Yes Template Acc
IQUAD [224] 75K 30 rooms AI2-THOR Multiple-choice Yes Template Acc
MT-EQA [225] 19.3K 588 envs House3D simulator Binary answer Yes Template Acc
MP3D-EQA [226] 1,136 83 envs MatterPort3D Answer set 53 Yes Template Acc
EgoVQA [227] 600 16 IU Multi-view Multiple-choice (1 out of 5) No Human annotators Acc
EgoTaskQA [228] 40K 2K LEMMA Answer set (open-answer and binary verifications) No Human annotators for relationship triples, then template Acc
EQA-MX [229] 8.2M 750K images CAESAR simulator Answer set No Question templates, answer set (simulator ground truth) Acc
OpenEQA [230] 557 + 1079 180 envs HM3D + ScanNet Open-answer (LLM scorer) No Human annotators Score
VI Challenges and Future Directions
Safety first
Safety is paramount in robotics, as robots interact directly with the physical world. Ensuring the safety of robotic systems requires the integration of real-world commonsense. This involves the incorporation of robust safety guardrails, risk assessment frameworks, and human-robot interaction protocols. RLHF and “evaluation without execution” can also significantly lower safety risks [37]. Interpretability and expandability of VLA decision-making processes are also crucial for enhancing robot safety through error diagnosis and troubleshooting.
Datasets & Benchmarks
In addition to the issues discussed in §V, comprehensive benchmarks that cover a wide range of skills, objects, embodiments, and environments remain to be developed. Moreover, metrics beyond the success rate are needed for a fine-grained diagnosis of issues in VLA models, as highlighted by EAI [222] for LLMs.
Foundation Models & Generalization
VLA foundation models or robotic foundation models (RFM) for embodied AI remains an open research topic primarily due to the diversity in embodiments, environments, and tasks. Many have made significant progress [139, 147], but still lacks generalization capability on par with LLMs in NLP. Attaining such a level of generalization is very challenging, as it requires the development of many core AGI capabilities.
Multi-Modality
VLAs inherit many challenges associated with multimodal models, such as obtaining useful embeddings and aligning different modalities. Current approaches, like ImageBind [231] and LanguageBind [232], align different modalities to the image or language embedding space, respectively. Within a unified embedding space, MLLMs can accommodate diverse modalities and become more general. However, whether focusing on embeddings alone is sufficient remains under debate. Although beyond the scope of VLAs, other modalities—such as audio [60] and haptics [233]—have proven useful for certain embodied AI applications. Incorporating additional modalities is often desirable, but they also introduce extra complexity into the model design.
Framework for Long-Horizon Tasks
The hierarchical framework is currently the most practical approach for long-horizon tasks. However, it increases system complexity and potential points of failure. Frequent task execution failures can trigger re-planning, which can cause significant latency. Moreover, monolithic task planners share similar architectures with LVLAs, so employing two large models can be redundant and may hinder scalability. Additionally, although modular task planners typically do not require training, they are not plug-and-play: language-based models may generate subtasks that control policies cannot execute, whereas code-based models require modules to be pre-wrapped in APIs manually. Therefore, developing a unified framework that directly translates long-horizon tasks into low-level control signals in an end-to-end fashion is worth exploring.
Real-Time Responsiveness
Unlike conversational AI, many robotic applications require real-time decision-making to respond to dynamic environments. If inference time cannot keep pace with environmental changes, the model may generate obsolete actions repeatedly. However, current VLA models—especially LVLAs and task planners—face a trade-off between speed and capacity. Novel mechanisms are thus needed to strike an optimal balance.
Multi-agent Systems
Cooperative multi-agent systems offer benefits such as distributed perception and collaborative fault recovery. However, they also face challenges, including effective communication, coordinated dispatching, and fleet heterogeneity. In certain scenarios, individual agents may have conflicting goals, which further increases complexity.
Ethical and Societal Implications
Robotics has always raised various ethical, legal, and societal concerns. These include risks related to privacy, safety, job displacement, bias in decision-making, and the impact on social norms and human relationships.
Applications
Most current VLAs focus on household and industrial settings [2], which shape the available datasets and models. Given the potential of VLA models, a wide range of new applications is possible. One particularly important application is in healthcare, such as surgical robots [234] and care robots [235]. Healthcare demands higher safety and privacy standards and may necessitate novel techniques such as human-in-the-loop (HITL) control and federated learning. Moreover, due to the significant domain gap, specialized vision models might be needed for medical images.
VII Conclusion
Vision-language-action models hold immense promise for enabling embodied agents to interact with the physical world and fulfill users’ instructions. This paper is the first survey to review large VLAs alongside other generalized VLAs. Our taxonomy provides a high-level overview of three main lines of research: key components, control policies, and task planners. We meticulously summarize and compare their model designs and technical details. Additionally, we highlight essential resources for training and evaluating VLAs, such as datasets, simulators, and benchmarks. We hope this survey captures the rapidly evolving landscape of embodied AI and inspires future research.